OpenStudy (anonymous):

Two positive numbers have a difference of 12. The larger number is 5 more than twice the smaller. Find the two numbers

OpenStudy (anonymous):

Try to express your conditions in equations with your unknow numbers beiing x and y.

OpenStudy (tad1):

x - y = 12. x = 5 + 2y. Solve the system of equations 5 +2y - y = 12 y = 7 so x = 5 +14 = 19 then ck 19-7 = 12 and x = 5 + 2(7)= 19